OpenStudy (anonymous):

Write the equation in logarithmic form. 2^-3=0.125

OpenStudy (anonymous):

I'm really confused. I don't understand logarithms at all

OpenStudy (anonymous):

@freckles ??

OpenStudy (freckles):

\[a^x=y \\ \log_a(y)=x\]

OpenStudy (freckles):

these two forms are equivalent

OpenStudy (freckles):

identify what you a, x, and y are and just plug them into the second equation I wrote

OpenStudy (anonymous):

Ohhh ok thanks.

OpenStudy (freckles):

Yeah I didn't read your question at first. :p Anyways here is an example if you need one: \[3^{2}=9 \\ \log_3(9)=2\]

OpenStudy (anonymous):

I got log2(1/8)=-3

OpenStudy (freckles):

superb victory!

OpenStudy (freckles):

\[\log_2(\frac{1}{8})=-3 \]

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (freckles):

good job you!